gene rides 10 minutes to a friends house and then walks 15 minutes to the gym. he rides 10 km/hr faster than he walks and the entire distance covered is 2.75 km. how far is it from the friends house to the gym?
5 Step process ?!
let [ ] statements....

Answer by ptaylor(2198)


Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
Let r=rate that he walks
Then r+10=rate that he rides
Distance covered riding=(r+10)*(1/6) {10 min=1/6 hr}
Distance covered walking=r*(1/4) (friend's house to gym)
Now we are told that the above two distance add up to 2.75 km, so:
(1/6)(r+10)+(1/4)r=2.75 multiply each term by 12
2(r+10)+3r=33
2r+20+3r=33 subtract 20 from each side
5r=33-20
5r=13
r=2.6 km/hr-------------rate that he walks
r+10=2.6+10=12.6 km/hr----rate that he rides
So
Distance from friend's house to gym=(2.6)*(1/4)=0.65 km
CK
(1/6)*(12.6)+(1/4)(2.6)=2.75
2.1+0.65=2.75
2.75=2.75
Hope this helps---ptaylor